# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v1

@@ C.1 Experiments with synthetic datasets
We employ a latent variable model (Bishop & Nasrabadi, 2006) for generating synthetic multimodal datasets. A latent variable model is a statistical model for data ${\bf X}\in\mathbb{R}^{d_{x}}$, under which ${\bf X}$ is generated according to a conditional probability distribution $P_{{\bf X}|{\bf Z}}$, where ${\bf Z}\in\mathbb{R}^{d_{z}}$ is the latent variable. In terms of the representation learning framework, ${\bf Z}$ can be seen as a true representation of ${\bf X}$. Moreover, we assume that the class label ${\bf Y}\in[K]$ and the latent variable ${\bf Z}$ are jointly distributed according to $P_{{\bf Z},{\bf Y}}$.
For the marginal distribution of ${\bf Z}$, we make use of a Gaussian mixture model (GMM) (Bishop & Nasrabadi, 2006), and hence the probability density function (PDF) of ${\bf Z}$ is a weighted sum of Gaussian densities. In particular, the PDF of ${\bf Z}$ is defined as follows:
@@
where ${\rm sigmoid}(x)=\frac{1}{1+e^{-x}}$ is applied element-wise, and $\Theta_{i}^{(1)}\in\mathbb{R}^{d_{x}\times d_{z}}$ and $\Theta_{i}^{(2)}\in\mathbb{R}^{d_{x}\times d_{x}}$ are matrices randomly generated from Gaussian distribution. Moreover, after $\Theta_{i}^{(1)},i\in[M]$ are generated, we set arbitrary columns of them all zero, so that the number of all zero columns decreases in $i$. For example, $60\%$ of columns of $\Theta_{1}^{(1)}$ are all-zero, while only $10\%$ of columns of $\Theta_{M}^{(1)}$ are all-zero. This enables approximate control the correlation between ${\bf X}_{i}$ and ${\bf Z}$, providing estimates of best modality (${\bf X}_{M}$) or worst modality (${\bf X}_{1}$). To have meaningful labels for this latent model, which requires for downstream tasks, we set the labels ${\bf Y}$ being the component index in GMM. In particular, since there are $K$ components in GMM (26), there exists $K$ categories in ${\bf Y}$. We conduct experiments with three different synthetic datasets by setting $M=4,6,8$. For all synthetic datasets, we fix $d_{x}=16$, $d_{z}=8$, and $K=50$.
@@
We initialize two different versions of backbones for all modalities, where the first is a random backbone (highlighted by (rnd) in figures), and the second is a backbone pretrained with InfoNCE loss. For each backbone, we use a simple multilayer perceptron (MLP). Comparing the results with these two versions of backbone provides how much both FABind and CentroBind are robust to backbone quality. Given the backbones for $M$ modalities, we align the corresponding embedding spaces using either FABind with anchor ${\bf X}_{i}$ (denoted by ${\bf X}_{i}$-B in figures) or CentroBind (denoted by CB in figures). Finally, with the encoders aligned by either FABind or CentroBind, we evaluate classification accuracy as a measure of representation quality. We use a simple MLP for the classifier. To distinguish between accuracy with embeddings from a single modality and the one with concatenated embeddings from all modalities, we denote by acc(${\bf Z}_{i}$) the accuracy with embeddings from $i$-th modality and by acc(All) the accuracy with embeddings from all modalities.
Additional experimental results on synthetic datasets with $M=6,8$ number of modalities are shown in Figure 3 and Figure 4.

### v2

@@ C.1 Experiments with synthetic datasets
We employ a latent variable model (Bishop & Nasrabadi, 2006) for generating synthetic multi-modal datasets. A latent variable model is a statistical model for data ${\bf X}\in\mathbb{R}^{d_{x}}$, under which ${\bf X}$ is generated according to a conditional probability distribution $P_{{\bf X}|{\bf Z}}$, where ${\bf Z}\in\mathbb{R}^{d_{z}}$ is the latent variable. In terms of the representation learning framework, ${\bf Z}$ can be seen as a low dimensional representation of ${\bf X}$. We assume that the class label ${\bf Y}\in[K]$ and the latent variable ${\bf Z}$ are jointly distributed according to $P_{{\bf Z},{\bf Y}}$. In our setting, we exploit Gaussian mixture model (GMM) (Bishop & Nasrabadi, 2006) for the latent variable ${\bf Z}$, and we generate $M$ modalities ${\bf X}_{i}=g_{i}({\bf Z})+{\bf N},i\in[M]$ with random noise ${\bf N}$ and some non-linear projections $g_{i}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{x}}$. We choose the projections in a way such that each model can be ranked in ascending order, i.e., ${\bf X}_{1}$ is the worst, and ${\bf X}_{4}$ is the best modality in terms of their inherent correlation with the latent variable. The class label ${\bf Y}$ is set to the component id of GMM.
In particular, the PDF of ${\bf Z}$ is defined as follows:
@@
where ${\rm sigmoid}(x)=\frac{1}{1+e^{-x}}$ is applied element-wise, and $\Theta_{i}^{(1)}\in\mathbb{R}^{d_{x}\times d_{z}}$ and $\Theta_{i}^{(2)}\in\mathbb{R}^{d_{x}\times d_{x}}$ are matrices randomly generated from Gaussian distribution. Moreover, after $\Theta_{i}^{(1)},i\in[M]$ are generated, we set arbitrary columns of them all zero, so that the number of all zero columns decreases in $i$. For example, $60\%$ of columns of $\Theta_{1}^{(1)}$ are all-zero, while only $10\%$ of columns of $\Theta_{M}^{(1)}$ are all-zero. This enables approximate control the correlation between ${\bf X}_{i}$ and ${\bf Z}$, providing estimates of best modality (${\bf X}_{M}$) or worst modality (${\bf X}_{1}$). To have meaningful labels for this latent model, which requires for downstream tasks, we set the labels ${\bf Y}$ being the component index in GMM. In particular, since there are $K$ components in GMM (27), there exist $K$ categories in ${\bf Y}$. We conduct experiments with three different synthetic datasets by setting $M=4,6,8$. For all synthetic datasets, we fix $d_{x}=16$, $d_{z}=8$, and $K=50$.
@@
We initialize two different versions of backbones for all modalities, where the first is a random backbone (highlighted by (rnd) in figures), and the second is a backbone pretrained with InfoNCE loss. For each backbone, we use a simple multilayer perceptron (MLP). Comparing the results with these two versions of backbone provides how much both FABind and CentroBind are robust to backbone quality. Given the backbones for $M$ modalities, we align the corresponding embedding spaces using either FABind with anchor ${\bf X}_{i}$ (denoted by ${\bf X}_{i}$-B in figures) or CentroBind (denoted by CB in figures). Finally, with the encoders aligned by either FABind or CentroBind, we evaluate classification accuracy as a measure of representation quality. We use a simple MLP for the classifier. To distinguish between accuracy with embeddings from a single modality and the one with concatenated embeddings from all modalities, we denote by acc(${\bf Z}_{i}$) the accuracy with embeddings from $i$-th modality and by acc(All) the accuracy with embeddings from all modalities. Specifically, for acc(All), we fuse the multi-modal embeddings using MLP layers. Therefore, the accuracy of the multi-modal case without binding methods (e.g., $\times$ method and the multi-modal column in Table 3) can be considered a naive baseline for multi-modal learning.
Comparison with baseline methods.
Figure 3 shows the validation accuracy of each method (without binding, FABind with anchor ${\bf X}_{1}$, FABind with anchor ${\bf X}_{4}$, and CentroBind). For the same experimental setting, Figure 5 includes additional accuracy curves for $acc({\bf Z}_{1})$ and $acc(All)$. For better readability, the corresponding accuracy is provided in Table 3.
We conduct experiments with two types of backbone encoders: randomly initialized backbones and pre-trained backbones. For each type, we extract embeddings using four different methods: representations without binding (denoted by $\times$ in Table 5), FABind with anchor modality ${\bf X}_{1}$ (denoted as FABind-${\bf X}_{1}$), FABind with anchor modality ${\bf X}_{4}$ (denoted as FABind-${\bf X}_{4}$), and CentroBind. The embedding quality is then evaluated using classification accuracy. Specifically, we train five different decoders for each case: four uni-modal decoders (one for each modality) and one multi-modal decoder for the concatenated embeddings of all modalities. The results show that CentroBind outperforms the other baseline methods. Notably, CentroBind demonstrates superior performance in the case of randomly initialized backbones, indicating robustness to poor backbone quality.
Additional experimental results on synthetic datasets with $M=6$ and $M=8$ modalities are presented in Figure 8 and Figure 9, respectively. These results exhibit similar trends to those observed with $M=4$ modalities. These experiments verify that CentroBind is capable of handling a large number of modalities effectively.
(a) Pre-trained backbones are used.
(b) Randomly initialized backbones are used.
Table 3: Classification accuracies presented in Figure 3.
Backbone Method uni-modal multi-modal ${\bf X}_{1}$ ${\bf X}_{2}$ ${\bf X}_{3}$ ${\bf X}_{4}$ ${\bf X}_{1},\cdots,{\bf X}_{4}$ Pre-trained $\times$ 0.2166 0.2878 0.3536 0.3923 0.6985 FABind-${\bf X}_{1}$ 0.2180 0.2736 0.3210 0.2999 0.5541 FABind-${\bf X}_{4}$ 0.2483 0.3349 0.4207 0.3896 0.7024 CentroBind 0.2540 0.3433 0.4162 0.4559 0.6974 Random $\times$ 0.2109 0.2472 0.2597 0.2815 0.6648 FABind-${\bf X}_{1}$ 0.2119 0.2587 0.3034 0.3081 0.5502 FABind-${\bf X}_{4}$ 0.2447 0.3076 0.3826 0.2813 0.6742 CentroBind 0.2582 0.3392 0.4224 0.4649 0.7006
(a) t-SNE: FABind.
(b) t-SNE: CentroBind.
(c) UMAP: FABind.
(d) UMAP: CentroBind.
Figure 6: Representation visualization via t-SNE and UMAP.
Representation visualization.
Figure 6 presents t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018) visualizations of embeddings generated by FABind and CentroBind. For this visualization, we use synthetic datasets with $4$ modalities, ensuring that each modality is equally informative, and plot the embeddings for ${\bf X}_{1}$. FABind is anchored at ${\bf X}_{4}$, and both binding methods utilize pre-trained backbones.
In both t-SNE and UMAP visualizations, CentroBind produces more clustered representations, whereas FABind results in more scattered embeddings. These findings validate our analysis that CentroBind creates a superior representation space by effectively learning both intra and shared information.
Figure 7: Training loss.
Convergence and stability analysis.
The convergence rate of CentroBind may differ from that of FABind due to the replacement of the fixed anchor with a dynamic anchor. In Figure 7, we plot the loss curves of CentroBind and FABind during training. The results show that the loss of CentroBind saturates earlier than that of FABind. We attribute this to the fact that the centroid serves as a minimizer of embeddings in terms of Euclidean distance, making it easier to converge embeddings to their centroid compared to converging them to one specific embedding.
The plot also reveals a crossover point where the loss curves intersect. We believe this occurs due to the number of InfoNCE losses optimized by CentroBind and FABind. Specifically, with $M$ modalities, CentroBind minimizes $M$ InfoNCE losses, while FABind minimizes $M-1$ InfoNCE losses. This results in a smaller loss for FABind when the encoders are well-trained, which explains the crossover point observed in Figure 7.
Table 4: Classification accuracies presented in Figure 10. In the experiment in Figure 10(a) and 10(b), ${\bf X}_{1},{\bf X}_{2}$, and ${\bf X}_{3}$ are very noisy, and ${\bf X}_{4}$ is highly informative. In the experiment in Figure 10(c) and 10(d), ${\bf X}_{1}$ and ${\bf X}_{2}$ are very noisy, and ${\bf X}_{3}$ and ${\bf X}_{4}$ are highly informative. We choose ${\bf X}_{4}$ for FABind for the best-fixed anchor modality for both cases. The weighted average method uses prior knowledge of modality quality to determine the weights for each modality. Random anchor method without intra-learning uses a randomly chosen modality as an anchor for each iteration under a fixed anchor encoder, while with intra-learning we train intra-modal learning by not freezing the anchor encoder.
Backbone Method Figure 10(a) and 10(b) Figure 10(c) and 10(d) ${\bf X}_{2}$ ${\bf X}_{4}$ ${\bf X}_{1},\cdots,{\bf X}_{4}$ ${\bf X}_{2}$ ${\bf X}_{4}$ ${\bf X}_{1},\cdots,{\bf X}_{4}$ Pre-trained $\times$ 0.115 0.296 0.566 0.099 0.256 0.537 FABind-${\bf X}_{4}$ 0.124 0.297 0.639 0.115 0.263 0.540 CentroBind 0.131 0.363 0.618 0.116 0.336 0.563 Weighted average 0.133 0.342 0.609 0.102 0.338 0.574 Random + intra learning 0.131 0.347 0.613 0.105 0.353 0.554 Random anchor 0.147 0.359 0.619 0.097 0.327 0.579 Median (coordinate-wise) 0.134 0.363 0.634 0.112 0.375 0.582 Random $\times$ 0.092 0.114 0.487 0.067 0.176 0.465 FABind-${\bf X}_{4}$ 0.131 0.143 0.575 0.112 0.153 0.523 CentroBind 0.132 0.355 0.626 0.113 0.336 0.559 Weighted average 0.115 0.347 0.619 0.109 0.336 0.556 Random + intra learning 0.132 0.333 0.602 0.104 0.309 0.562 Random anchor 0.145 0.354 0.618 0.097 0.324 0.552 Median (coordinate-wise) 0.137 0.347 0.612 0.112 0.330 0.565
Comparison with other adaptive anchor generation.
We compare the centroid-based adaptive anchor method with other potential approaches, such as weighted average, random anchor fixing, and component-wise median. Figure 10 illustrates the accuracies of each method under scenarios where modalities are unevenly distributed. Specifically, we create $4$ modalities with differing quality levels. In experiments (a) and (b) of Figure 10, ${\bf X}_{1}$, ${\bf X}_{2}$, and ${\bf X}_{3}$ are set as highly uninformative, while ${\bf X}_{4}$ represents a high-quality dataset. Conversely, experiments (c) and (d) use ${\bf X}_{1}$ and ${\bf X}_{2}$ as poor-quality datasets, while ${\bf X}_{3}$ and ${\bf X}_{4}$ are high-quality datasets.
For the weighted average method (denoted as WAB in Figure 10), we assign weights based on modality quality: $(0.2,0.2,0.2,1)$ for experiments (a) and (b), and $(0.2,0.2,0.8,0.8)$ for experiments (c) and (d). These weights correspond to the information rate of each modality.
For the random modality dynamic anchor method (denoted as RB in Figure 10), we randomly select one modality as the dynamic anchor at each iteration, with the anchor encoder frozen. To investigate the impact of intra-modal learning, we also conduct experiments with a random anchor that includes intra information learning (denoted as RB+Intra). In this case, the anchor modality is randomly selected at each iteration, and the anchor encoder is not frozen, allowing all encoders to be trained.
Since the median is more robust to outliers than the average (Lopuhaä & Rousseeuw, 1991), we additionally evaluate the case of a median-based dynamic anchor. In high-dimensional spaces, rather than in the univariate case, a coordinate-wise median can be used as a naive generalization of the univariate median to the multivariate setting, preserving its robustness to outliers. We assess the dynamic anchor binding method using the coordinate-wise median approach (denoted as MB in Figure 10). Specifically, for the median anchor, we compute the $j$th coordinate of the $i$th anchor as ${\boldsymbol{a}}_{i,j}={\rm Median}({\boldsymbol{z}}_{1,i,j},{\boldsymbol{z}}_{2,i,j},\cdots,{\boldsymbol{z}}_{M,i,j})$, where ${\boldsymbol{z}}_{m,i,j}$ denotes the $j$th coordinate of the embedding for the $i$th sample in modality $m$. For improved readability, we summarize the final accuracies for each method and modality in Table 4.
This scenario, where modal distributions are uneven, is commonly referred to as the modality imbalance problem (Du et al., 2023; Peng et al., 2022; Zhang et al., 2024). Intuitively, in the presence of modality imbalance, the centroid may produce suboptimal dynamic anchor constructions, and other methods, such as weighted averages, might yield better results. Nevertheless, CentroBind consistently performs better or comparably to weighted average methods, demonstrating its robustness to the modality imbalance problem.
From these experiments, we conjecture that the specific dynamic anchor generation method may not significantly impact final performance, provided that all encoders are well-trained during the process.
Addressing the modality imbalance problem typically requires additional information, such as domain knowledge, labels, or downstream task insights. Since this work focuses on multi-modal alignment under contrastive learning, we do not assume such information is available. We therefore leave the exploration of the modality imbalance problem for dynamic anchor generation as a direction for future work.
Table 5: Classification accuracy evaluated on each modality (training and evaluation modalities are the same) with MUStARD dataset. Asterisk denotes different backbone encoders and pretraining settings.
[…]
@@
(a) Pre-trained encoders are used.
(b) Random backbones are used.
(c) Pre-trained backbones are used.
(d) Random backbones are used.
Figure 10: Comparison of other dynamic anchor generation methods. (a) and (b): Modal qualities are set to $(0.2,0.2,0.2,1)$. (c) and (d): Modal qualities are set to $(0.2,0.2,0.8,0.8)$. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind method to backbones with anchor ${\bf X}_{i}$ or applying CentroBind; WAB: weighted average for dynamic anchor with weight identical to the predefined quality for each modality; RB+Intra: randomly choosing a modality for a dynamic anchor in every iteration and intra information learning; RB: randomly choosing a modality for a dynamic anchor in every iteration; MB: coordinate-wise median for dynamic anchors; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (ran): if random backbones are used.
@@ C.2 Experiments with real-world datasets
UniBind
We evaluate UniBind as a baseline method, using LLM-generated descriptions as the anchor modality. Specifically, UniBind generates descriptions for each modality using a large language model (LLM), ensuring that every modality is paired with corresponding descriptions. These descriptions collectively form a knowledge base, and UniBind optimizes the InfoNCE loss between each modality and its paired description from the knowledge base. In this framework, the anchor modality is the LLM-augmented representation. It is important to note that the LLM-generated descriptions for different modality pairs can vary, which may hinder effective multi-modal alignment (see Table 2). In our experiments, we generate descriptions for video and audio modalities using the VideoLLaMA2.1-7B-AV audio-visual model from VideoLLaMA2 (Cheng et al., 2024), and for the text modality, we use the Qwen2.5-32B-Instruct model from Qwen2.5 (Team, 2024). We evaluate UniBind’s performance in two settings: standard classification accuracy (Table 5) and zero-shot cross-modal classification (Table 2).
AudioCLIP
We employ AudioCLIP (Guzhov et al., 2022), which aligns image, text, and audio representations into a unified multi-modal space. To extend its capabilities to the video modality in our experiments, we adapt AudioCLIP to extract embeddings for video, audio, and text modalities using a pretrained model. For audio, we follow AudioCLIP’s approach, padding audio samples to ensure uniform input sizes. For text, we utilize its pretrained settings, truncating tokenized text to 77 tokens, which only occurs in one instance. For the video modality, we use the center frame as a representative image sample. Finally, embeddings from all three modalities are concatenated for downstream tasks.
ViT-Lens
In our experiments, we leverage the pretrained models from ViT-Lens to extract embeddings for audio, text, and video modalities. We generally follow the example code provided by the authors. [^6] https://github.com/TencentARC/ViT-Lens Note that we select the center frame image from the video to extract the embedding.
Classification results.
In contrast to the cross-modal retrieval results in Table 1 and zero-shot cross-modal classification in Table 2, Table 5 presents the classification accuracy of FABind, UniBind, and CentroBind for each modality as well as for multi-modal scenarios. Specifically, embeddings are extracted using the binding methods, and a simple decoder is trained to classify the embeddings. In Table 5, we report the sarcasm and speaker classification accuracies of decoders trained and evaluated on the same modality.
For sarcasm detection, CentroBind generally outperforms other baseline methods. While UniBind performs poorly in cross-modal classification, it achieves better performance in speaker classification compared to others. This improvement is due to the LLM-augmented descriptions, which provide additional knowledge (from LLMs) to the embeddings. Notably, UniBind utilizes $4$ modalities, whereas FABind and CentroBind only use $3$, which could penalize the performance of FABind and CentroBind . Nevertheless, CentroBind consistently outperforms FABind. Moreover, our method can also incorporate LLM-augmented descriptions as an additional modality, potentially improving its performance further.
Although a direct comparison is not feasible, we also include the sarcasm detection accuracy of FactorCL (Liang et al., 2024b), SimMMDG (Dong et al., 2023), AudioCLIP (Guzhov et al., 2022), and ViT-Lens (Lei et al., 2024) for reference. ViT-Lens, in particular, achieves higher performance than CentroBind due to its use of larger backbone encoders, such as Vision Transformer (ViT) (Khan et al., 2022) and pretraining on extremely large-scale datasets. However, since ViT-Lens can be considered a variant of FABind, applying our dynamic anchor method could further improve its performance. Specifically, ViT-Lens uses a pretrained CLIP model as the anchor encoder, while the other non-anchored modalities use pretrained ViT models with modality adaptation layers. Within our framework, CentroBind could adopt the pretrained Vision Transformer as backbone encoders, potentially enhancing its performance further.
C.3 Experiments with pre-trained backbone
Figure 11: Top-1 retrieval accuracy of ImageBind and CentroBind on DreamBooth datset.
In addition to experiments on synthetic and real-world datasets, we evaluate our method using a pre-trained ImageBind backbone for image and text modalities. We utilize the DreamBooth dataset (Ruiz et al., 2023), which comprises 30 subjects across 15 different classes of image-text, to fine-tune the pre-trained encoders. The evaluation is conducted through a retrieval task, effectively measuring the quality of the learned unified embedding space.
Specifically, we begin by loading the pre-trained ImageBind model (Girdhar et al., 2023) and fine-tuning its text and image encoders on the DreamBooth training dataset using both the ImageBind and CentroBind methods. For fine-tuning, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 4, resulting in 5.4 million trainable parameters out of a total of 1.2 billion parameters. After fine-tuning, we measure the top-1 retrieval accuracy on the test dataset. The optimization process uses the AdamW optimizer with the following hyperparameters: a batch size of 16, a learning rate of $5.0\times 10^{-6}$, a maximum of 500 epochs, a weight decay of $10^{-4}$, and a temperature parameter set to 0.07. In this experiment, we omit the text augmentation step because the text samples in the DreamBooth dataset are too short. Instead, we create a centroid-based dynamic anchor using the embeddings of augmented images and original text.
The results presented in Figure 11 demonstrate that CentroBind outperforms the ImageBind method, even when leveraging a powerful pre-trained backbone and the image modality. We anticipate that this performance gap would widen if the backbone were less powerful or if the image modality were excluded. These findings corroborate our analysis of the dynamic anchor binding method, highlighting its effectiveness in enhancing multi-modal representations.